OpenStudy (anonymous):

Shawn and Ellie are construction workers. Shawn can complete a concrete job in 4 hours, while Ellie can complete it in 6 hours. The foreman says that it will take them 5 hours to complete it if they work together. Explain each step in solving this equation, and determine if the foreman is correct or not.

OpenStudy (anonymous):

So far I got... Shawn: 1/4 Ellie: 1/6 1/4 + 1/6 = 1/x

OpenStudy (dtan5457):

Shawn: 1/4 Ellie: 1/6 1/4 + 1/6 = 1/x that part is good so far, so now you want to eliminate the fractions with I'd say 12x. 3x+2x=12 now that's simply 5x=12 which is 2.4 hours the foreman is wrong